# Simple Lifelong Learning Machines

### v21

@@ I Introduction
Figure 1: Performance summary on different vision and speech benchmark datasets. Columns are different evaluation criteria (see Section III for definitions, and Section VI for experimental details), each strip of colored dots corresponds to an algorithm (we introduce SiLLy-N here) and each dot represents a task. Older tasks have darker colors. Resource growing algorithms have a ‘*’. EWC, O-EWC, SI, TAG and ER always perform worse than LwF, and hence we do not show them in the plot. SiLLy-N (red) outperforms all reference algorithms in terms of overall (first column), forward (second column), and backward (third column). Importantly, such better transfer is achieved at high overall accuracy (fourth column). For CORe50 dataset, we are unable to run Model Zoo in the resource constrained environment of the experiment and LwF completely forgets the first several tasks with the task-accuracies going down to the chance level ($20\%$). See Appendix Figure 1 for the error bars for each dot in the figure.
@@ III-B Lifelong learning evaluation criteria
Definition 2 (Forward Transfer)
@@
Definition 3 (Backward Transfer)
@@ III-C Computational Taxonomy of Lifelong Learners
($n,T$) ($n,T$) ($n\propto T$) $(n,T)$ $(n\propto T)$
parametric $1$ $1$ $1$ $1$ $1$ SiLLy-N-M, SiLLy-F-M
parametric $1$ $1$ $1$ $n$ $n$ O-EWC [12], SI [14], LwF [15]
parametric $1$ $T$ $n$ $nT$ $n^{2}$ EWC [11]
parametric $1$ $n$ $n$ $nT$ $n^{2}$ Total Replay
@@ IV-A Our representation ensembling algorithms
Figure 3: Space and time complexity as a function on number of tasks in CIFAR 10X10. Section VI describes the experimental setup and architecture of the convolutional network used. Top left: Model size, SiLLy can flexibly operate between resource constrained and growing modes. Top right: Memory consumed by SiLLy-N is dominated by the encoder size. Bottom left: SiLLy takes less training time compared to other baselines. Bottom right: Inference time taken by different lifelong learners for $1000$ samples. Bottom row has log y-axis.
@@ IV-A3 Additional realization of our approach using random forest as encoder
Figure 3 top left shows the model size for our proposed approach grows linearly with the number of tasks. Moreover, the memory consumed by the new channels is negligible compared to the memory required to store the encoders (Figure 3 top right). The time required for inference on $1000$ testing points (Figure 3 bottom right) is an order of magnitude lower compared to the time required to train a new encoder with $500$ samples(Figure 3 bottom left).
@@ V Simulation Data Study
Figure 4: SiLLy-N demonstrates forward and backward transfer. (A) 750 samples from: (Ai) Gaussian XOR, (Aii) XNOR, which has the same optimal discriminant boundary as XOR, and (Aiii) R-XOR, which has a discriminant boundary that is uninformative, and therefore adversarial, to XOR. (Bi) Generalization error for XOR, and (Bii) XNOR. SiLLy-N outperforms DN on XOR when XNOR data is available, and on XNOR when XOR data are available. (Biii) Forward and backward transfer of SiLLy-N are positive for all sample sizes. (Ci) In an adversarial task setting, SiLLy-N gracefully forgets XOR, whereas DN  catastrophically forget and interfere. (Cii) $\mathsf{Backward~{}Transfer}$ is maximum positive with respect to XOR when the optimal decision boundary of $\theta$-XOR is similar to that of XOR (e.g. angles far from $45^{\circ}$), and negative otherwise. The dashed line shows the regression line fitted through the original points. (Ciii) $\mathsf{Backward~{}Transfer}$ is a nonlinear function of the source training sample size (XOR sample size is fixed at $500$).
@@ V-B Forward and backward transfer for adversarial tasks
To further investigate this relationship, we design a suite of R-XOR examples, generalizing R-XOR from only $45^{\circ}$ to any rotation angle between $0^{\circ}$ and $90^{\circ}$, sampling 100 points from XOR, and another 100 from each R-XOR (Figure 4Cii). Note that we could not run the experiment for a lot of Monte Carlo repetition to have a smooth curve and hence we have shown a regressed curve fitted to the low repetition noisy curve. As the angle increases from $0^{\circ}$ to $45^{\circ}$, $\mathsf{Backward~{}Transfer}$ gradually decreases for SiLLy-N. The $45^{\circ}$-XOR is the maximally adversarial R-XOR. Thus, as the angle further increases, $\mathsf{Backward~{}Transfer}$ increases back up to $\approx 0.18$ at $90^{\circ}$, which has an identical discriminant boundary to XOR. Moreover, when $\theta$ is fixed at $25^{\circ}$, $\mathsf{Backward~{}Transfer}$ increases at different rates for different sample sizes of the source task (Figure 4Ciii).
@@ VI-A Reference algorithms
We compared our approaches to $16$ reference lifelong learning methods. Among them five are resource growing as well as modular approach: ProgNN [21], DF-CNN [22], LMC [25], Model Zoo[20], CoSCL [61]. Note that ”Model Zoo” was published after our work was archived on arXiv, and the authors have built on our work (personal communications). Other reference algorithms are resource constrained: Elastic Weight Consolidation (EWC) [11], Online-EWC (O-EWC) [12], Synaptic Intelligence (SI) [14], Learning without Forgetting (LwF) [15], RanDumb [62] and “None”.
@@ VI-B3 FOOD1k 50X20 Dataset
In this experiment, we use Food1k which is a large scale vision dataset consisting of $1000$ food categories from Food2k [69]. FOOD1k 50X20 splits these data into $50$ tasks with $20$ classes each. For each class, we randomly sampled $60$ samples per class for training the models and used rest of the data for testing purpose. Because on the CIFAR experiments Model Zoo performs the best among the reference resource growing models, and LwF is the best performing resource constrained algorithm, we only use them as the reference models for the large scale experiment to avoid heavy computational cost. As shown in Figure 1 third column, SiLLy-N performs the best among all the algorithms on this large dataset.
@@ VI-C Ablation experiments
In this experiment, we ablate the capability of SiLLy-N to grow additional resources after learning $4$ encoders. We also reduce the number of channels and nodes at each encoder layer by four times to keep the total number of parameters similar to the other constant-resource-algorithms. As shown in the top row of Figure 5, SiLLy-N-4 still shows positive forward and backward transfer with constant resources. However, the accuracy for SiLLy-N-4 gets reduced compared to that of resource growing SiLLy-N in Figure 1. Note that all the baseline algorithms have negative backward transfer. This experiment indicates that constant resource mode operation for SiLLy-N may be advantageous when we have a lot of tasks to learn and have a decent amount of storage budget available. We will elaborate the above point later with a large scale dataset (food1k).
@@
In this experiment, we train four different versions of SiLLy-N sequentially on the $10$ tasks from CIFAR 10X10. The only difference between different versions of the algorithms is the amount of old task data replayed. In four different versions of each algorithm, we replay $40\%$, $60\%$, $80\%$ and $100\%$ of the old task data respectively. As apparent from Figure 5 bottom, replaying old task data has no effect on forward transfer, but replaying more data improves backward transfer as the number of tasks increases.
@@
We explore the effect of using pretrained encoders on the performance of SiLLy-N. For this experiment only, we use ResNet 50 and vision transformer ViT_B16 (provided in keras-vit package) [70]. We freeze all the layers excluding the final two linear layers during training. Pretraining the encoders results in better accuracy and forward transfer, but less backward transfer for SiLLy-N (Figure 6). Vision transformers achieved lower accuracy, as expected with the small samples sizes used here [71]. Pretrained network requires less training epochs for each task (with early stopping). In this experiment, one forward pass through ViT takes $\sim 5$ seconds and training a ViT encoder ($20$ epochs) takes $\sim 4$ minutes using an Apple M1 Max chip and $64$ GB of RAM.
@@ VI-D Adversarial analysis
We repeat the same experiment mentioned above $5$ times more by permuting the task order. Figure 7 left column shows the mean as well as the spread of $\mathsf{Transfer}$ over all tasks remain the same for different shuffled task orders. Note that the encoders and thus all the channels remain the same regardless of the above task order permutation after all the tasks have been introduced, hence the distribution of $\mathsf{Transfer}$ for all the tasks after all the tasks have been introduced remain similar.
@@ VI-E Constant Resource Mode Operation
Figure 8: SiLLy-N in constant resource mode operation. Improvement in Transfer between tasks becomes negligible at nearly $30$ encoders. The user can choose to operate in lower budget with less transfer.
@@ VI-E1 FOOD1k 50X20
The binary distinction we made above, algorithms either build resources or reallocate them, is a false dichotomy, and biologically unnatural. In biological learning, systems develop from building to fixed resources, as they grow from juveniles to adults. To explore this continuum of amount of resources to grow, we experiment on FOOD1k 50X20 dataset using the constant resource mode operation of SiLLy-N as described in Section IV. We evaluate the performance of SiLLy-N for different number of encoder budget. Performance of SiLLy-N saturates after $30$ encoders, though with only $5$ encoders, still demonstrates forward and backward transfer (Figure 8).
@@ VI-E2 CORe50 110X5
To further evaluate the effectiveness of constant resource mode operation, we utilize a large-scale dataset: CORe50 [72], partitioned into $110$ tasks with $5$ classes each and $100$ training samples per class. To simulate a resource-constrained environment, we conducted all experiments on an Apple M1 Max chip with 64 GB of RAM. We set a budget of 20 encoders, as this configuration maintains high accuracy (above $90\%$) across all tasks. The fourth row of Figure 1 shows that SiLLy-N achieves higher $\mathsf{Transfer}$ operating in constant resource mode on a large dataset, that is, CORe50 compared to the smaller datasets discussed above. In particular, $\mathsf{BackwardTransfer}$ drops to zero after 20 tasks due to the encoder limit, leaving only $\mathsf{ForwardTransfer}$ for subsequent tasks. Moreover, we cannot run resource-growing algorithms like Model Zoo in this resource-constrained environment. The constant resource algorithm, such as LwF, completely forgets the first several tasks, and the corresponding task accuracies go as low as the chance level, $20\%$.
@@ VI-F SiLLy-F on tabular data
In this experiment, we experiment with SiLLy-F, an additional realization of our approach using random forests as encoders (described in Section IV). We flatten the CIFAR 10X10 data and use them as tabular data. We train two other best performing baseline algorithms, SiLLy-N and Model Zoo and use three fully connected hidden layers, each having $2000$ nodes, as encoders. As shown in Figure 9, SiLLy-F performs the best among all the approaches. This experiment shows our approach can be used as a general structure to do lifelong learning using other machine learning models as encoder.
@@ A-A3 Training Procedures
7:function SiLLy-X.fit($t,(\mathbf{x}^{t},\mathbf{y}^{t})$)
8:     $u_{t},\mathcal{I}_{OOB}^{t}\leftarrow$ encoder.fit($\mathbf{x}^{t}$, $\mathbf{y}^{t}$) $\triangleright$ train an encoder on training data partitioned into in-bag and OOB samples
@@
6:$v_{t}$triangleright channel for task $t$
7:function SiLLy-X.add_channel($t,u_{t},(\mathbf{x}_{t},\mathbf{y}_{t}),\mathcal{I}_{OOB}^{t}$)
8:     $v_{t}\leftarrow u_{t}$.add_channel($(\mathbf{x}_{t},\mathbf{y}_{t}),\mathcal{I}_{OOB}^{t}$) $\triangleright$ add the new in-task channel using OOB data
@@
10:         $v_{t}\leftarrow u_{t^{\prime}}$.update_channel($\mathbf{x}_{t},\mathbf{y}_{t},v_{t}$) $\triangleright$ update the channel for task $t$ using the old encoders
@@
6:$\mathcal{V}=\{\boldsymbol{v}_{t^{\prime}}\}_{t^{\prime}=1}^{t-1}$
7:function SiLLy-X.update_old_channel($t,u_{t},\mathcal{D},\mathcal{V}$)
@@
9:         $v_{t^{\prime}}\leftarrow u_{t}$.update_channel($\mathcal{D}^{t^{\prime}},v_{t^{\prime}}$) $\triangleright$ update the old task channels
@@
6:$\hat{y}$triangleright a predicted class label
7:function $\hat{y}=$ SiLLy-X.predict($t,\boldsymbol{x},v_{t}$)
@@
max_samples (OOB split) $0.67$
n_estimators (decision forest channel) $20$
@@
n_estimators $10$
max_depth $30$
max_samples (OOB split) $0.67$
min_samples_leaf $1$
@@ A-B Reference Algorithm Implementation Details
The same network architecture was used for all baseline deep learning methods. Following the work in  [30], the ‘base network architecture’ consisted of five convolutional layers followed by two-fully connected layers each containing 2000 nodes with ReLU non-linearities and a softmax output layer. The convolutional layers had $16$, $32$, $64$, $128$ and $254$ channels, they used batch-norm and a ReLU non-linearity, they had a 3x3 kernel, a padding of 1 and a stride of 2 (except the first layer, which had a stride of 1). This architecture was used with a multi-headed output layer (i.e., a different output layer for each task) for all algorithms using a fixed-size network. For ProgNN and DF-CNN the same architecture was used for each column introduced for each new task, and in our SiLLy-N this architecture was used for the transformers $u_{t}$ (see above). In these implementations, ProgNN and DF-CNN have the same architecture for each column introduced for each task. Among the reference algorithms, EWC, O-EWC, LwF, SI, Total Replay and Partial Replay results were produced using the repository https://github.com/GMvandeVen/progressive-learning-pytorch. For ProgNN and DF-CNN  we used the code provided in https://github.com/Lifelong-ML/DF-CNN. For all other reference algorithms, we modified the code provided by the authors to match the deep net architecture as mentioned above and used the default hyperparameters provided in the code.
@@ A-C Training Time Complexity Analysis
We use the soft-O notation $\tilde{\mathcal{O}}$ to quantify complexity [75]. Letting $n$ be the sample size and $T$ be the number of tasks, we write that the capacity, space or time complexity of a lifelong learning algorithm is $f(n,t)=\tilde{\mathcal{O}}(g(n,T))$ when $|f|$ is bounded above asymptotically by a function $g$ of $n$ and $T$ up to a constant factor and polylogarithmic terms. For simplifying the calculation, we make the following assumptions:
@@
Assumption $3$ enables us to write time complexity as a function of the sample size. Table I summarizes the capacity, space and time complexity of several reference algorithms, as well as our SiLLy-N and SiLLy-F.
@@ A-D Simulation Experiment Details
In each simulation, we constructed an environment with two tasks. For each, we sample 750 times from the first task, followed by 750 times from the second task. These 1,500 samples comprise the training data. We sample another 1,000 hold out samples to evaluate the algorithms. For SiLLy-N, we have used a deep network (DN) architecture with two hidden layers each having $10$ nodes. Similarly, for SiLLy-N experiments we did $100$ repetitions and reported the results after smoothing it using moving average with a window size of $5$.

### v22

@@ I Introduction
[…]
@@ III-B Lifelong learning evaluation criteria
Definition 2 (Forward Transfer).
@@
Definition 3 (Backward Transfer).
@@ III-C Computational Taxonomy of Lifelong Learners
[…]
parametric 111 111 111 111 111 SiLLy-N-M, SiLLy-F-M
parametric 111 111 111 $n$ $n$ O-EWC [12], SI [14], LwF [15]
parametric 111 $T$ $n$ $nT$ $n^{2}$ EWC [11]
parametric 111 $n$ $n$ $nT$ $n^{2}$ Total Replay
@@ IV-A Our representation ensembling algorithms
Figure 3: Space and time complexity as a function on number of tasks in CIFAR 10X10. Section VI describes the experimental setup and architecture of the convolutional network used. Top left: Model size, SiLLy can flexibly operate between resource constrained and growing modes. Top right: Memory consumed by SiLLy-N is dominated by the encoder size. Bottom left: SiLLy takes less training time compared to other baselines. Bottom right: Inference time taken by different lifelong learners for 100010001000 samples. Bottom row has log y-axis.
@@ IV-A3 Additional realization of our approach using random forest as encoder
Figure 3 top left shows the model size for our proposed approach grows linearly with the number of tasks. Moreover, the memory consumed by the new channels is negligible compared to the memory required to store the encoders (Figure 3 top right). The time required for inference on 100010001000 testing points (Figure 3 bottom right) is an order of magnitude lower compared to the time required to train a new encoder with 500500500 samples(Figure 3 bottom left).
@@ V Simulation Data Study
Figure 4: SiLLy-N demonstrates forward and backward transfer. (A) 750 samples from: (Ai) Gaussian XOR, (Aii) XNOR, which has the same optimal discriminant boundary as XOR, and (Aiii) R-XOR, which has a discriminant boundary that is uninformative, and therefore adversarial, to XOR. (Bi) Generalization error for XOR, and (Bii) XNOR. SiLLy-N outperforms DN on XOR when XNOR data is available, and on XNOR when XOR data are available. (Biii) Forward and backward transfer of SiLLy-N are positive for all sample sizes. (Ci) In an adversarial task setting, SiLLy-N gracefully forgets XOR, whereas DN  catastrophically forget and interfere. (Cii) $\mathsf{Backward~Transfer}$ is maximum positive with respect to XOR when the optimal decision boundary of $\theta$-XOR is similar to that of XOR (e.g. angles far from $45^{\circ}$), and negative otherwise. The dashed line shows the regression line fitted through the original points. (Ciii) $\mathsf{Backward~Transfer}$ is a nonlinear function of the source training sample size (XOR sample size is fixed at 500500500).
@@ V-B Forward and backward transfer for adversarial tasks
To further investigate this relationship, we design a suite of R-XOR examples, generalizing R-XOR from only $45^{\circ}$ to any rotation angle between $0^{\circ}$ and $90^{\circ}$, sampling 100 points from XOR, and another 100 from each R-XOR (Figure 4Cii). Note that we could not run the experiment for a lot of Monte Carlo repetition to have a smooth curve and hence we have shown a regressed curve fitted to the low repetition noisy curve. As the angle increases from $0^{\circ}$ to $45^{\circ}$, $\mathsf{Backward~Transfer}$ gradually decreases for SiLLy-N. The $45^{\circ}$-XOR is the maximally adversarial R-XOR. Thus, as the angle further increases, $\mathsf{Backward~Transfer}$ increases back up to 0.18$\approx$ 0.18≈ 0.18 at $90^{\circ}$, which has an identical discriminant boundary to XOR. Moreover, when $\theta$ is fixed at $25^{\circ}$, $\mathsf{Backward~Transfer}$ increases at different rates for different sample sizes of the source task (Figure 4Ciii).
@@ VI-A Reference algorithms
We compared our approaches to 161616 reference lifelong learning methods. Among them five are resource growing as well as modular approach: ProgNN [22], DF-CNN [23], LMC [26], Model Zoo[21], CoSCL [63]. Note that ”Model Zoo” was published after our work was archived on arXiv, and the authors have built on our work (personal communications). Other reference algorithms are resource constrained: Elastic Weight Consolidation (EWC) [11], Online-EWC (O-EWC) [12], Synaptic Intelligence (SI) [14], Learning without Forgetting (LwF) [15], RanDumb [64] and “None”.
@@ VI-B3 FOOD1k 50X20 Dataset
In this experiment, we use Food1k which is a large scale vision dataset consisting of 100010001000 food categories from Food2k [71]. FOOD1k 50X20 splits these data into 505050 tasks with 202020 classes each. For each class, we randomly sampled 606060 samples per class for training the models and used rest of the data for testing purpose. Because on the CIFAR experiments Model Zoo performs the best among the reference resource growing models, and LwF is the best performing resource constrained algorithm, we only use them as the reference models for the large scale experiment to avoid heavy computational cost. As shown in Figure 1 third column, SiLLy-N performs the best among all the algorithms on this large dataset.
@@ VI-C Ablation experiments
In this experiment, we ablate the capability of SiLLy-N to grow additional resources after learning 444 encoders. We also reduce the number of channels and nodes at each encoder layer by four times to keep the total number of parameters similar to the other constant-resource-algorithms. As shown in the top row of Figure 5, SiLLy-N-4 still shows positive forward and backward transfer with constant resources. However, the accuracy for SiLLy-N-4 gets reduced compared to that of resource growing SiLLy-N in Figure 1. Note that all the baseline algorithms have negative backward transfer. This experiment indicates that constant resource mode operation for SiLLy-N may be advantageous when we have a lot of tasks to learn and have a decent amount of storage budget available. We will elaborate the above point later with a large scale dataset (food1k).
@@
[…]
@@
We explore the effect of using pretrained encoders on the performance of SiLLy-N. For this experiment only, we use ResNet 50 and vision transformer ViT_B16 (provided in keras-vit package) [72]. We freeze all the layers excluding the final two linear layers during training. Pretraining the encoders results in better accuracy and forward transfer, but less backward transfer for SiLLy-N (Figure 6). Vision transformers achieved lower accuracy, as expected with the small samples sizes used here [73]. Pretrained network requires less training epochs for each task (with early stopping). In this experiment, one forward pass through ViT takes 5$\sim$ 5∼ 5 seconds and training a ViT encoder (202020 epochs) takes 4$\sim$ 4∼ 4 minutes using an Apple M1 Max chip and 646464 GB of RAM.
@@ VI-D Adversarial analysis
We repeat the same experiment mentioned above 555 times more by permuting the task order. Figure 7 left column shows the mean as well as the spread of $\mathsf{Transfer}$ over all tasks remain the same for different shuffled task orders. Note that the encoders and thus all the channels remain the same regardless of the above task order permutation after all the tasks have been introduced, hence the distribution of $\mathsf{Transfer}$ for all the tasks after all the tasks have been introduced remain similar.
@@ VI-E Constant Resource Mode Operation
Figure 8: SiLLy-N in constant resource mode operation. Improvement in Transfer between tasks becomes negligible at nearly 303030 encoders. The user can choose to operate in lower budget with less transfer.
@@ VI-E1 FOOD1k 50X20
The binary distinction we made above, algorithms either build resources or reallocate them, is a false dichotomy, and biologically unnatural. In biological learning, systems develop from building to fixed resources, as they grow from juveniles to adults. To explore this continuum of amount of resources to grow, we experiment on FOOD1k 50X20 dataset using the constant resource mode operation of SiLLy-N as described in Section IV. We evaluate the performance of SiLLy-N for different number of encoder budget. Performance of SiLLy-N saturates after 303030 encoders, though with only 555 encoders, still demonstrates forward and backward transfer (Figure 8).
@@ VI-E2 CORe50 110X5
[…]
@@ VI-F SiLLy-F on tabular data
In this experiment, we experiment with SiLLy-F, an additional realization of our approach using random forests as encoders (described in Section IV). We flatten the CIFAR 10X10 data and use them as tabular data. We train two other best performing baseline algorithms, SiLLy-N and Model Zoo and use three fully connected hidden layers, each having 200020002000 nodes, as encoders. As shown in Figure 9, SiLLy-F performs the best among all the approaches. This experiment shows our approach can be used as a general structure to do lifelong learning using other machine learning models as encoder.
@@ A-A3 Training Procedures
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
max_samples (OOB split) 0.670.670.67
n_estimators (decision forest channel) 202020
@@
n_estimators 101010
max_depth 303030
max_samples (OOB split) 0.670.670.67
min_samples_leaf 111
@@ A-B Reference Algorithm Implementation Details
The same network architecture was used for all baseline deep learning methods. Following the work in  [31], the ‘base network architecture’ consisted of five convolutional layers followed by two-fully connected layers each containing 2000 nodes with ReLU non-linearities and a softmax output layer. The convolutional layers had 161616, 323232, 646464, 128128128 and 254254254 channels, they used batch-norm and a ReLU non-linearity, they had a 3x3 kernel, a padding of 1 and a stride of 2 (except the first layer, which had a stride of 1). This architecture was used with a multi-headed output layer (i.e., a different output layer for each task) for all algorithms using a fixed-size network. For ProgNN and DF-CNN the same architecture was used for each column introduced for each new task, and in our SiLLy-N this architecture was used for the transformers $u_{t}$ (see above). In these implementations, ProgNN and DF-CNN have the same architecture for each column introduced for each task. Among the reference algorithms, EWC, O-EWC, LwF, SI, Total Replay and Partial Replay results were produced using the repository https://github.com/GMvandeVen/progressive-learning-pytorch. For ProgNN and DF-CNN  we used the code provided in https://github.com/Lifelong-ML/DF-CNN. For all other reference algorithms, we modified the code provided by the authors to match the deep net architecture as mentioned above and used the default hyperparameters provided in the code.
@@ A-C Training Time Complexity Analysis
[…]
@@
Assumption 333 enables us to write time complexity as a function of the sample size. Table I summarizes the capacity, space and time complexity of several reference algorithms, as well as our SiLLy-N and SiLLy-F.
@@ A-D Simulation Experiment Details
In each simulation, we constructed an environment with two tasks. For each, we sample 750 times from the first task, followed by 750 times from the second task. These 1,500 samples comprise the training data. We sample another 1,000 hold out samples to evaluate the algorithms. For SiLLy-N, we have used a deep network (DN) architecture with two hidden layers each having 101010 nodes. Similarly, for SiLLy-N experiments we did 100100100 repetitions and reported the results after smoothing it using moving average with a window size of 555.